OpenStudy (anonymous):

Ciara solved the exponential equation 3x+1 = 15 and her work is shown below. What is the first step she did incorrectly? Step 1: log 3x+1 = log15 Step 2: (x + 1)log 3 = log15 Step 3: log3 = log 15 over x plus 1 Step 4: 0.477121 = 1.176091 over x plus 1 Step 5: 0.477121(x + 1) = 1.176091 Step 6: x + 1 = 1.176091 over 0.477121 Step 7: x + 1 = 2.464975 Step 8: x = 1.464975

OpenStudy (solomonzelman):

The equation: \(\color{#000000 }{ \displaystyle 3^{x+1}=15 }\) Step 1: \(\color{#000000 }{ \displaystyle 3^{x+1}=15 }\) Step 2: \(\color{#000000 }{ \displaystyle \log(3^{x+1})=\log(15) }\) Step 3: \(\color{#000000 }{ \displaystyle \log(3)=\frac{\log(15)}{x+1} }\) Step 4: \(\color{#000000 }{ \displaystyle 0.477121 =\frac{1.176091}{ x+1} }\) Step 5: \(\color{#000000 }{ \displaystyle 0.477121(x+1) =1.176091 }\) Step 6: \(\color{#000000 }{ \displaystyle x+1 =\frac{1.176091}{ 0.477121} }\) Step 7: \(\color{#000000 }{ \displaystyle x + 1 = 2.464975 }\) Step 8: \(\color{#000000 }{ \displaystyle x = 1.464975 }\)

OpenStudy (solomonzelman):

That is the equation and those are the steps. Am I interpreting the question correctly?

OpenStudy (anonymous):

yep

OpenStudy (solomonzelman):

I don't think any of the steps were incorrectly done, however, some steps were a lot of extra work....

OpenStudy (solomonzelman):

my bad, step 1: \(\color{#000000 }{ \displaystyle\log(3^{x+1})=\log(15) }\) step 2: \(\color{#000000 }{ \displaystyle(x+1)\log(3)=\log(15) }\)

OpenStudy (solomonzelman):

And then all you need is: \(\color{#000000 }{ \displaystyle x+1=\frac{\log(15)}{\log(3)} }\)

OpenStudy (solomonzelman):

and then you are two steps from the solution.

OpenStudy (anonymous):

so im guessing step 3 is incorrect

OpenStudy (solomonzelman):

Yes. I would rather say that step 3 is a waste.

OpenStudy (anonymous):

ok... ty

OpenStudy (solomonzelman):

yw